OpenStudy (anonymous):

Find the equation of the parabola with focus (0,2) and directrix y = -2.

OpenStudy (anonymous):

\[4py=x^2\] with directrix x = -p substitute in and you'll have what you seek

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

oops, i mean directrix y = -p

OpenStudy (anonymous):

its okay, thank you for your help

OpenStudy (anonymous):

which form of the equation do you need?

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

standard form:\[ y = ax^2+bx+c\] vertex form, with vertex @ (h, k):\[y=a(x-h)^2+k\]

OpenStudy (anonymous):

i needed vertex form thank you for the formula

OpenStudy (anonymous):

do you know the vertex?